The planes pie1 and pie2 have cartesian equations x-y+z+1=0 and 2x+y-z-1=0 respectively.

(a) Show that these two planes are not parallel.

(b) Find a parametric equation of their line of intersection

(c)Give a cartesian equation of plane pie3 passing through point A(1,-1,-1) and perpendicular to pie1 and pie2.


1
Expert's answer
2021-06-08T14:31:36-0400

a) two planes "A_1x+B_1y+C_1z+D_1=0" and "A_2x+B_2y+C_2z+D_2=0" are parallel if "\dfrac{A_1}{A_2} = \dfrac{B_1}{B_2} = \dfrac{C_1}{C_2} ."

In our case "\dfrac{A_1}{A_2} = \dfrac{1}{2}, \; \dfrac{B_1}{B_2} = \dfrac{-1}{1}, \; \dfrac{C_1}{C_2} = \dfrac{1}{-1}" , these numbers are not the same, so the planes are not parallel.


b) the coordinates of points on the line of intersection satisfy the equations

"\begin{cases}
x-y+z+1 = 0, \\
2x+y-z-1 = 0
\end{cases}"

"\begin{cases}
x-y+z=-1, \\
2x+y-z=1
\end{cases}"

The equivalent matrices:

"\begin{bmatrix}
 1 & -1 & 1 & -1 \\
 2 & 1 & -1 & 1
\end{bmatrix}" , "\begin{bmatrix}
 1 & -1 & 1 & -1 \\
 0 & -\dfrac32 & \dfrac32 & -\dfrac32
\end{bmatrix}, 
\begin{bmatrix}
 1 & 0 & 0 & 0 \\
 0 & -\dfrac32 & \dfrac32 & -\dfrac32
\end{bmatrix},"

"\begin{bmatrix}
 1 & 0 & 0 & 0 \\
 0 & 1 & -1 &1
\end{bmatrix},"

"\begin{cases}
x = 0 + 0t, \\
y = 1 + t,\\
0 = 0 + t.
\end{cases}"


c) Let the equation of the third plane be "A_3x+B_3y+C_3z + D_3 = 0."

It is perpendicular to pi_1 and pi_2, so

"1\cdot A_3 - 1\cdot B_3 + 1\cdot C_3 = 0, \\
2\cdot A_3 + 1\cdot B_3 - 1\cdot C_3 = 0" , therefore "A_3 = 0," "B_3 = C_3." Let "B_3 = 1."

A(1,-1,-1) belongs to plane, so "A_3\cdot 1 + B_3\cdot (-1) + C_3\cdot(-1) + D_3 = 0,"

"0 - 1 - 1 + D_3 = 0, D_3 = 2."

So the equation of the plane is "y+z+2 = 0."